# Colorful Pinball: Density-Weighted Quantile Regression for Conditional Guarantee of Conformal Prediction

### v2

@@ A.5 Detailed results
Split 0.0071 ± 0.0021 0.0150 ± 0.0016 0.0054 ± 0.0009 0.0392 ± 0.0119 0.0065 ± 0.0010 0.0090 ± 0.0009 0.0108 ± 0.0024 0.0237 ± 0.0024
PLCP (G=20) 0.0044 ± 0.0013 0.0043 ± 0.0051 0.0023 ± 0.0012 0.0290 ± 0.0130 0.0043 ± 0.0023 0.0045 ± 0.0019 0.0055 ± 0.0042 0.0057 ± 0.0039
PLCP (G=50) 0.0041 ± 0.0010 0.0026 ± 0.0017 0.0021 ± 0.0009 0.0302 ± 0.0123 0.0037 ± 0.0023 0.0038 ± 0.0016 0.0073 ± 0.0036 0.0050 ± 0.0029
Gaussian-Scoring 0.0034 ± 0.0011 0.0019 ± 0.0005 0.0014 ± 0.0003 0.0287 ± 0.0094 0.0040 ± 0.0044 0.0060 ± 0.0013 0.0049 ± 0.0076 0.6455 ± 0.0073
CQR 0.0036 ± 0.0014 0.0037 ± 0.0010 0.0016 ± 0.0006 0.0284 ± 0.0168 0.0023 ± 0.0007 0.0027 ± 0.0007 0.0034 ± 0.0016 0.0116 ± 0.0010
CQR-ALD 0.0033 ± 0.0010 0.0020 ± 0.0007 0.0015 ± 0.0005 0.0311 ± 0.0138 0.0021 ± 0.0013 0.0025 ± 0.0007 0.0038 ± 0.0020 0.0100 ± 0.0014
RCP 0.0035 ± 0.0011 0.0046 ± 0.0007 0.0018 ± 0.0003 0.0087 ± 0.0030 0.0023 ± 0.0003 0.0031 ± 0.0009 0.0024 ± 0.0006 0.0069 ± 0.0012
RCP-ALD 0.0030 ± 0.0006 0.0028 ± 0.0006 0.0016 ± 0.0005 0.0138 ± 0.0077 0.0021 ± 0.0005 0.0028 ± 0.0007 0.0016 ± 0.0003 0.0058 ± 0.0009
CPCP 0.0032 ± 0.0009 0.0023 ± 0.0007 0.0016 ± 0.0005 0.0068 ± 0.0018 0.0010 ± 0.0003 0.0027 ± 0.0007 0.0029 ± 0.0009 0.0044 ± 0.0008
CPCP (Mix) 0.0028 ± 0.0007 0.0018 ± 0.0008 0.0016 ± 0.0003 0.0073 ± 0.0017 0.0010 ± 0.0002 0.0028 ± 0.0009 0.0023 ± 0.0012 0.0038 ± 0.0010
CPCP (Clip) 0.0029 ± 0.0005 0.0018 ± 0.0003 0.0015 ± 0.0004 0.0064 ± 0.0023 0.0011 ± 0.0002 0.0031 ± 0.0008 0.0020 ± 0.0008 0.0036 ± 0.0009
CPCP (Clip+Mix) 0.0026 ± 0.0006 0.0014 ± 0.0003 0.0015 ± 0.0003 0.0059 ± 0.0019 0.0010 ± 0.0002 0.0025 ± 0.0007 0.0014 ± 0.0004 0.0034 ± 0.0007
@@ B.3.1 Preliminary: fast rates via local Rademacher complexity
Without additional curvature assumptions, global complexity bounds typically yield the slow rate $\mathcal{E}_{q}(n)=O_{\mathbb{P}}(n^{-1/4})$. To obtain fast rates, we invoke the theory of local Rademacher complexity [bartlett2005local].
@@ B.3.4 Bounding the weight estimation error (Term II)
Hence, we can applying Bernstein’s inequality [bach2024learning] to transform the empirical norm into population norm, with probability at least $1-\zeta$:
@@ C.1 Gap between $F_{S\mid X}(\hat{q}(x))$ and $\pi(x)$
In this part, we supplement the results proved in plassier2025rectifying.
@@
Here, we adopt a slightly stronger  [^3] We assume the existence of PDF, while in  plassier2025rectifying, the continuity of CDF is assumed. version than plassier2025rectifying and suppress complex notations to enhance the readability. Assumption C.1 essentially characterizes the quality of quantile estimator, i.e., to what extent the quantile estimator removes the covariate-dependency on scores.
Theorem C.2 (Theorem 2 in plassier2025rectifying).

### v3

@@ A.5 Detailed results
Split 0.0602 ± 0.0071 0.1223 ± 0.0050 0.0374 ± 0.0034 0.1536 ± 0.0092 0.1044 ± 0.0046 0.0956 ± 0.0058 0.1120 ± 0.0058 0.1079 ± 0.0031
PLCP (G=20) 0.0424 ± 0.0167 0.0642 ± 0.0310 0.0198 ± 0.0075 0.0789 ± 0.0279 0.0809 ± 0.0295 0.0466 ± 0.0092 0.0573 ± 0.0262 0.0656 ± 0.0182
PLCP (G=50) 0.0351 ± 0.0142 0.0445 ± 0.0241 0.0219 ± 0.0075 0.0685 ± 0.0296 0.0557 ± 0.0331 0.0455 ± 0.0100 0.0417 ± 0.0102 0.0565 ± 0.0109
Gaussian-Scoring 0.0337 ± 0.0086 0.0330 ± 0.0057 0.0168 ± 0.0079 0.0826 ± 0.0327 0.0733 ± 0.0349 0.0443 ± 0.0105 0.0371 ± 0.0146 0.0691 ± 0.0120
CQR 0.0307 ± 0.0136 0.0384 ± 0.0082 0.0170 ± 0.0070 0.0783 ± 0.0282 0.0506 ± 0.0112 0.0270 ± 0.0099 0.0510 ± 0.0117 0.0762 ± 0.0035
CQR-ALD 0.0274 ± 0.0114 0.0301 ± 0.0131 0.0177 ± 0.0041 0.0476 ± 0.0268 0.0404 ± 0.0204 0.0328 ± 0.0127 0.0461 ± 0.0092 0.0776 ± 0.0051
RCP 0.0208 ± 0.0076 0.0446 ± 0.0083 0.0155 ± 0.0067 0.0353 ± 0.0184 0.0371 ± 0.0099 0.0329 ± 0.0108 0.0342 ± 0.0078 0.0574 ± 0.0052
RCP-ALD 0.0274 ± 0.0142 0.0373 ± 0.0100 0.0157 ± 0.0051 0.0557 ± 0.0343 0.0427 ± 0.0185 0.0309 ± 0.0084 0.0322 ± 0.0132 0.0495 ± 0.0059
RCP-MultiHead 0.0239 ± 0.0072 0.0484 ± 0.0082 0.0179 ± 0.0077 0.0366 ± 0.0138 0.0477 ± 0.0064 0.0323 ± 0.0084 0.0335 ± 0.0112 0.0573 ± 0.0070
CPCP 0.0195 ± 0.0097 0.0379 ± 0.0105 0.0129 ± 0.0047 0.0271 ± 0.0148 0.0166 ± 0.0062 0.0370 ± 0.0120 0.0295 ± 0.0076 0.0632 ± 0.0090
CPCP (Clip) 0.0228 ± 0.0103 0.0256 ± 0.0065 0.0148 ± 0.0066 0.0278 ± 0.0100 0.0189 ± 0.0065 0.0346 ± 0.0059 0.0242 ± 0.0070 0.0530 ± 0.0063
CPCP (Mix) 0.0196 ± 0.0083 0.0275 ± 0.0094 0.0135 ± 0.0048 0.0182 ± 0.0107 0.0158 ± 0.0054 0.0319 ± 0.0092 0.0219 ± 0.0070 0.0498 ± 0.0090
CPCP (Clip+Mix) 0.0178 ± 0.0076 0.0219 ± 0.0048 0.0116 ± 0.0067 0.0268 ± 0.0102 0.0155 ± 0.0061 0.0267 ± 0.0065 0.0182 ± 0.0050 0.0433 ± 0.0048
CPCP-$0.01$ 0.0225 ± 0.0082 0.0394 ± 0.0108 0.0170 ± 0.0070 0.0280 ± 0.0139 0.0153 ± 0.0057 0.0362 ± 0.0113 0.0308 ± 0.0092 0.0568 ± 0.0082
CPCP-$0.01$ (Clip+Mix) 0.0165 ± 0.0078 0.0230 ± 0.0081 0.0105 ± 0.0051 0.0202 ± 0.0113 0.0150 ± 0.0073 0.0257 ± 0.0065 0.0177 ± 0.0045 0.0464 ± 0.0065
CPCP-$0.05$ 0.0217 ± 0.0081 0.0375 ± 0.0129 0.0140 ± 0.0070 0.0250 ± 0.0094 0.0165 ± 0.0064 0.0337 ± 0.0081 0.0312 ± 0.0079 0.0575 ± 0.0125
CPCP-$0.05$ (Clip+Mix) 0.0169 ± 0.0082 0.0243 ± 0.0049 0.0110 ± 0.0053 0.0221 ± 0.0128 0.0152 ± 0.0051 0.0273 ± 0.0061 0.0175 ± 0.0056 0.0376 ± 0.0049
Table 9: Full comparison of $\ell_{2}$-ERT. $\ell_{2}$-ERT ($\downarrow$) on all benchmarks with all baselines.
Split 0.0060 ± 0.0015 0.0287 ± 0.0029 0.0044 ± 0.0008 0.0338 ± 0.0067 0.0219 ± 0.0028 0.0108 ± 0.0018 0.0163 ± 0.0020 0.0225 ± 0.0018
PLCP (G=20) 0.0030 ± 0.0023 0.0076 ± 0.0087 0.0011 ± 0.0011 0.0139 ± 0.0104 0.0139 ± 0.0087 0.0035 ± 0.0016 0.0059 ± 0.0051 0.0085 ± 0.0057
PLCP (G=50) 0.0022 ± 0.0019 0.0038 ± 0.0062 0.0011 ± 0.0014 0.0095 ± 0.0071 0.0077 ± 0.0094 0.0039 ± 0.0026 0.0031 ± 0.0014 0.0055 ± 0.0023
Gaussian-Scoring 0.0018 ± 0.0009 0.0024 ± 0.0009 0.0005 ± 0.0005 0.0127 ± 0.0130 0.0136 ± 0.0145 0.0024 ± 0.0012 0.0028 ± 0.0033 0.0079 ± 0.0035
CQR 0.0016 ± 0.0013 0.0024 ± 0.0009 0.0005 ± 0.0003 0.0104 ± 0.0069 0.0044 ± 0.0023 0.0007 ± 0.0014 0.0048 ± 0.0032 0.0084 ± 0.0011
CQR-ALD 0.0011 ± 0.0010 0.0016 ± 0.0017 0.0006 ± 0.0004 0.0049 ± 0.0075 0.0032 ± 0.0035 0.0013 ± 0.0016 0.0040 ± 0.0027 0.0125 ± 0.0025
RCP 0.0006 ± 0.0006 0.0029 ± 0.0011 0.0004 ± 0.0004 0.0023 ± 0.0020 0.0016 ± 0.0009 0.0014 ± 0.0015 0.0018 ± 0.0008 0.0048 ± 0.0011
RCP-ALD 0.0013 ± 0.0016 0.0022 ± 0.0011 0.0005 ± 0.0004 0.0072 ± 0.0110 0.0023 ± 0.0027 0.0012 ± 0.0010 0.0023 ± 0.0026 0.0039 ± 0.0013
RCP-MultiHead 0.0009 ± 0.0006 0.0036 ± 0.0011 0.0005 ± 0.0005 0.0019 ± 0.0012 0.0026 ± 0.0007 0.0012 ± 0.0009 0.0018 ± 0.0009 0.0045 ± 0.0016
CPCP 0.0006 ± 0.0006 0.0022 ± 0.0012 0.0003 ± 0.0002 0.0011 ± 0.0009 0.0004 ± 0.0003 0.0026 ± 0.0029 0.0015 ± 0.0007 0.0060 ± 0.0017
CPCP (Clip) 0.0008 ± 0.0006 0.0011 ± 0.0006 0.0003 ± 0.0003 0.0011 ± 0.0006 0.0005 ± 0.0003 0.0014 ± 0.0007 0.0010 ± 0.0005 0.0042 ± 0.0013
CPCP (Mix) 0.0005 ± 0.0005 0.0013 ± 0.0008 0.0003 ± 0.0002 0.0006 ± 0.0005 0.0003 ± 0.0002 0.0018 ± 0.0027 0.0008 ± 0.0005 0.0033 ± 0.0016
CPCP (Clip+Mix) 0.0004 ± 0.0005 0.0007 ± 0.0003 0.0002 ± 0.0002 0.0012 ± 0.0010 0.0003 ± 0.0003 0.0006 ± 0.0007 0.0005 ± 0.0003 0.0026 ± 0.0008
CPCP-$0.01$ 0.0007 ± 0.0006 0.0024 ± 0.0013 0.0005 ± 0.0004 0.0014 ± 0.0010 0.0003 ± 0.0003 0.0020 ± 0.0016 0.0015 ± 0.0008 0.0053 ± 0.0014
CPCP-$0.01$ (Clip+Mix) 0.0003 ± 0.0004 0.0009 ± 0.0007 0.0002 ± 0.0002 0.0007 ± 0.0005 0.0003 ± 0.0003 0.0005 ± 0.0008 0.0005 ± 0.0002 0.0033 ± 0.0012
CPCP-$0.05$ 0.0006 ± 0.0005 0.0022 ± 0.0017 0.0003 ± 0.0002 0.0009 ± 0.0006 0.0004 ± 0.0002 0.0018 ± 0.0014 0.0018 ± 0.0009 0.0053 ± 0.0027
CPCP-$0.05$ (Clip+Mix) 0.0003 ± 0.0005 0.0010 ± 0.0004 0.0002 ± 0.0002 0.0009 ± 0.0008 0.0003 ± 0.0002 0.0004 ± 0.0007 0.0005 ± 0.0003 0.0019 ± 0.0009
Table 10: Full comparison of Volume. Volume ($\downarrow$) on all benchmarks with all baselines.
Split 0.7661 ± 0.0222 0.3793 ± 0.0158 0.1794 ± 0.0202 -0.2741 ± 0.2199 -1.7916 ± 0.0618 0.9917 ± 0.0408 -1.0367 ± 0.0727 1.3004 ± 0.0207
PLCP (G=20) 0.7528 ± 0.0303 0.3520 ± 0.0149 0.1468 ± 0.0253 -0.7786 ± 0.2926 -1.8795 ± 0.1089 0.8815 ± 0.0325 -1.3365 ± 0.1637 1.0509 ± 0.1180
PLCP (G=50) 0.7419 ± 0.0571 0.3427 ± 0.0147 0.1529 ± 0.0305 -0.7513 ± 0.3553 -1.9906 ± 0.1200 0.8848 ± 0.0349 -1.4444 ± 0.0921 1.0260 ± 0.0843
Gaussian-Scoring -0.7455 ± 0.0498 -1.6312 ± 0.0612 0.0454 ± 0.0159 1.1504 ± 0.6506 1.1303 ± 1.5991 -0.6082 ± 0.0486 -1.1826 ± 0.7100 1.6395 ± 0.2775
CQR 0.8118 ± 0.0441 0.2990 ± 0.0084 0.0603 ± 0.0181 0.0665 ± 0.3057 -2.3191 ± 0.0515 0.8934 ± 0.0372 -2.4395 ± 0.1315 0.6992 ± 0.0189
CQR-ALD 1.0122 ± 0.0889 0.3170 ± 0.0344 0.0524 ± 0.0187 0.6268 ± 0.2757 -2.4332 ± 0.0748 0.9521 ± 0.0495 -2.2226 ± 0.1615 0.9541 ± 0.0243
RCP 0.7402 ± 0.0638 0.3197 ± 0.0126 0.1580 ± 0.0288 -0.9301 ± 0.2656 -2.0543 ± 0.0667 0.9009 ± 0.0842 -1.5827 ± 0.0699 0.7489 ± 0.0249
RCP-ALD 0.7380 ± 0.0415 0.3219 ± 0.0125 0.1406 ± 0.0227 -0.8320 ± 0.3071 -2.0520 ± 0.0795 0.8611 ± 0.0318 -1.5375 ± 0.0824 0.7571 ± 0.0292
RCP-MultiHead 0.7339 ± 0.0269 0.3156 ± 0.0080 0.1546 ± 0.0242 -1.0368 ± 0.2576 -2.0542 ± 0.0608 0.9002 ± 0.0713 -1.5889 ± 0.0894 0.7492 ± 0.0258
CPCP 0.7423 ± 0.0468 0.3723 ± 0.0420 0.1456 ± 0.0386 -0.9894 ± 0.3102 -2.0635 ± 0.0782 0.9247 ± 0.1059 -1.4728 ± 0.1361 0.9208 ± 0.0541
CPCP (Clip) 0.7512 ± 0.0532 0.3459 ± 0.0137 0.1555 ± 0.0334 -0.8944 ± 0.3215 -2.0523 ± 0.0568 0.8767 ± 0.0333 -1.5072 ± 0.0840 0.8472 ± 0.0457
CPCP (Mix) 0.7370 ± 0.0433 0.3566 ± 0.0317 0.1380 ± 0.0308 -1.0269 ± 0.2389 -2.0764 ± 0.0482 0.9123 ± 0.0669 -1.5034 ± 0.1004 0.8262 ± 0.0469
CPCP (Clip+Mix) 0.7486 ± 0.0330 0.3381 ± 0.0120 0.1511 ± 0.0423 -0.9101 ± 0.1810 -2.0793 ± 0.0402 0.8778 ± 0.0236 -1.5591 ± 0.0597 0.7697 ± 0.0461
CPCP-$0.01$ 0.7483 ± 0.0492 0.3928 ± 0.1133 0.1582 ± 0.0478 -0.9517 ± 0.2808 -2.0613 ± 0.0403 0.9197 ± 0.0769 -1.4551 ± 0.0713 0.9045 ± 0.0804
CPCP-$0.01$ (Clip+Mix) 0.7553 ± 0.0373 0.3437 ± 0.0132 0.1501 ± 0.0358 -0.9692 ± 0.2858 -2.0506 ± 0.0754 0.8669 ± 0.0286 -1.5704 ± 0.0712 0.7614 ± 0.0477
CPCP-$0.05$ 0.7471 ± 0.0400 0.4128 ± 0.1902 0.1439 ± 0.0343 -1.0389 ± 0.2635 -2.0583 ± 0.0536 0.8895 ± 0.0424 -1.4253 ± 0.1310 0.8521 ± 0.1029
CPCP-$0.05$ (Clip+Mix) 0.7473 ± 0.0445 0.3340 ± 0.0118 0.1347 ± 0.0244 -1.0265 ± 0.2652 -2.0449 ± 0.0639 0.8724 ± 0.0329 -1.5644 ± 0.0611 0.7037 ± 0.0349
@@ B.3.1 Preliminary: fast rates via local Rademacher complexity
Without additional curvature assumptions, global complexity bounds typically yield the slow rate $\mathcal{E}_{q}(n)=O_{\mathbb{P}}(n^{-1/4})$. To obtain fast rates, we invoke the theory of local Rademacher complexity [Bartlett et al., 2005].
@@ B.3.4 Bounding the weight estimation error (Term II)
Hence, we can applying Bernstein’s inequality [Bach, 2024] to transform the empirical norm into population norm, with probability at least $1-\zeta$:
@@ C.1 Gap between $F_{S\mid X}(\hat{q}(x))$ and $\pi(x)$
In this part, we supplement the results proved in Plassier et al. [2025a].
@@
Here, we adopt a slightly stronger  [^2] We assume the existence of PDF, while in  Plassier et al. [2025a], the continuity of CDF is assumed. version than Plassier et al. [2025a] and suppress complex notations to enhance the readability. Assumption C.1 essentially characterizes the quality of quantile estimator, i.e., to what extent the quantile estimator removes the covariate-dependency on scores.
Theorem C.2 (Theorem 2 in Plassier et al. [2025a]).